OpenStudy (anonymous):

ntegrate sin(2x) dx why is the answer -1/2 cos(2x)+2 why the 1/2

OpenStudy (anonymous):

yeah your answer is right and the 1/2 is there so that when you differentiate your answer you can get back to sin(2x) and its negative because the derivative of cos(x)=-sin(x)

OpenStudy (anonymous):

why 1/2?

OpenStudy (anonymous):

is there like a formula?

OpenStudy (lgbasallote):

are you familiar with the u-substitution? because i'm going to use it...if you don't know it just tell me and i can solve it by the basic method ;) personally i prefer not to because it confuses me lol anyway... \[\int \sin (2x) dx\] let u = 2x \(\rightarrow\) du = 2dx \(\rightarrow \frac{du}{2} = dx\) \[\int \sin (2x)dx \implies \int \sin (u) \frac{du}{2}\] \[\implies \frac12 \int \sin u du\] integral of sin u is -cos u therefore \[\frac 12 \int \sin u du = -\frac{1}{2} \cos u\] but dont forget to sub back \[-\frac{1}{2} \cos u \implies -\frac 12 \cos (2x)\]

OpenStudy (lgbasallote):

if you're not familiar with u-sub...i'll use the good old messy method \[\int \sin (2x) dx\] this is not a perfect integral you need a 2dx for it to be a perfect integral \[\int [\sin (2x)] (2dx)\] now it's a perfect integral..but you need to multiply by 1/2 because you multiplied by 2 (so that nothing will change in terms of value) \[\frac{1}{2} \int \sin (2x) (2dx)\] the integral of that is just \(-\cos (2x)\) so... \[\frac 12 \int \sin (2x) (2dx) \implies -\frac 12 \cos (2x)\] u-sub is easier to understand isnt it. lol

OpenStudy (lgbasallote):

do you get either of the methods i wrote @wilbert06 ?

OpenStudy (anonymous):

@igbasallote I dont quite fully understand them maybe if you would do an examample I could see it

OpenStudy (anonymous):

or u actually did

OpenStudy (anonymous):

yes i do thank you

OpenStudy (lgbasallote):

for future references...it's Lgbasallote not igbasallote ;) and glad to help ^_^

OpenStudy (anonymous):

haha thank you!